OpenStudy (anonymous):

hi plz solve this: a+b=10 b+c=20 c+a= 24 a+b+c= ?

OpenStudy (anonymous):

Given details - a+b=10 b+c=20 a+c=24 and find a+b+c = ? Firstly sum all three equations 2a + 2b + 2c = 54 Divide by 2 a + b + c = 54/2 a + b + c = 27

OpenStudy (anonymous):

hope this helps :)

OpenStudy (anonymous):

medal me

OpenStudy (texaschic101):

a + b = 10 a = 10 - b b + c = 20 c = 20 - b a + c = 24 10 - b + 20 - b = 24 30 - 2b = 24 -2b = 24 - 30 -2b = -6 b = 3 a + b = 10 a + 3 = 10 a = 10 - 3 a = 7 b + c = 20 3 + c = 20 c = 20 - 3 c = 17 check.. a + b = 10 7 + 3 = 10 10 = 10 (correct) b + c = 20 3 + 17 = 20 20 = 20 (correct) a + c = 24 7 + 17 = 24 24 = 24 (correct) so a = 7, b = 3, and c = 17 a + b + c = 7 + 3 + 17 = 27

OpenStudy (anonymous):

@Liza14 thanks for your efford. your answer is less time consuming. But i need clerification.